OpenStudy (anonymous):

if a^3 - a -1 =0 how to solve a^4 +a^3 - a^2 -2a +9?

OpenStudy (anonymous):

a^3=a+1 a^4=a^3.a=a(a+1)=a^2+a a^4 +a^3 - a^2 -2a +9=a^2+a+a+1-a^2-2a+9=10